# Assemble Your Crew: Automatic Multi-agent Communication Topology Design via Autoregressive Graph Generation

### v2

@@ Abstract
Multi-agent systems (MAS) based on large language models (LLMs) have emerged as a powerful solution for dealing with complex problems across diverse domains. The effectiveness of MAS is critically dependent on its collaboration topology, which has become a focal point for automated design research. However, existing approaches are fundamentally constrained by their reliance on a template graph modification paradigm with a predefined set of agents and hard-coded interaction structures, significantly limiting their adaptability to task-specific requirements. To address these limitations, we reframe MAS design as a conditional autoregressive graph generation task, where both the system composition and structure are designed jointly. We propose ARG-Designer, a novel autoregressive model that operationalizes this paradigm by constructing the collaboration graph from scratch. Conditioned on a natural language task query, ARG-Designer sequentially and dynamically determines the required number of agents, selects their appropriate roles from an extensible pool, and establishes the optimal communication links between them. This generative approach creates a customized topology in a flexible and extensible manner, precisely tailored to the unique demands of different tasks. Extensive experiments across six diverse benchmarks demonstrate that ARG-Designer not only achieves state-of-the-art performance but also enjoys significantly greater token efficiency and enhanced extensibility. The source code of ARG-Designer is available at https://github.com/Shiy-Li/ARG-Designer.
Preprint. Under review.
@@ Introduction
Agents built on large language models (LLMs) have demonstrated impressive capabilities in tackling complex tasks across diverse domains, including code generation, data analysis, decision-making, and question answering (Zhu et al. 2024; Li et al. 2024a; Song et al. 2023; Wang et al. 2024; Zhong, Wang, and Shang 2024). To overcome the limitations of a single agent in tackling more complex tasks, the research interests have increasingly shifted towards multi-agent systems (MAS), which unlock new potential through collaborative interactions among agents with diverse capabilities and roles. Central to MAS is its collaboration topology, a graph that defines how agents with various roles are structured and how they exchange information. A growing body of evidence, spanning from sequential reasoning pipelines to debate-based approaches, demonstrates that MAS performance varies dramatically depending on how inter-agent communication is architected (Zhang et al. 2024; Zhou et al. 2025). Therefore, designing an effective collaboration graph tailored to specific tasks becomes a critical research challenge.
@@
Early research on MAS topology design focused on static and manually designed graphs, such as chains that enforce a sequential workflow (Wei et al. 2022; Hong et al. 2023), trees that enable structured deliberation (Yao et al. 2023), and fully connected graphs that ensure sufficient communication. Although these canonical collaboration topologies can facilitate effective coordination in specific scenarios, the inherent rigidity of these fixed topologies limits their adaptability across diverse tasks, often resulting in sub-optimal performance. To enhance flexibility and efficiency, a more recent line of work focuses on adaptively constructing task-specific communication structures using graph learning models (Zhuge et al. 2024; Shen et al. 2025). For example, AgentPrune (Zhang et al. 2025a) and AgentDropout (Wang et al. 2025) learn to create sparse and task-specific graphs by pruning connections or agents from a predefined template topology. More advanced approaches like G-Designer (Zhang et al. 2025b) follow the paradigm of graph structure learning, which leverages a graph autoencoder to learn efficient collaboration structures in a task-adaptive manner.
Despite their varied designs, existing graph learning-based methods often follow a shared paradigm: template graph modification (Fig. 1a). That is, they typically start from a fixed communication template based on a predefined set of agents and hard-coded interaction structures, and apply learnable adjustments, such as edge reweighting or pruning, to adapt the topology to specific tasks (Zhang et al. 2025a, b). Despite offering reasonable adaptability in constrained settings, this paradigm exhibits two inherent limitations. Limitation 1: Redundant Composition. To ensure structural flexibility, template graphs are often initialized with numerous agent roles and densely connected edges, many of which are unnecessary for a specific task. Even with pruning mechanisms, irrelevant agents or connections may be retained in the learned task-specific topology, leading not only to reduced efficiency but also to potential sub-optimal decision-making during execution. Limitation 2: Limited Extensibility. In the fast-evolving field of LLM-based agents, a massive number of new agent functionalities are emerging with increasing frequency. However, trained on a fixed template graph, the existing methods struggle to generalize to scenarios with dynamic agent sets or evolving collaboration needs. Meanwhile, it would be prohibitively expensive to build a large-scale template graph that covers all possible agent roles and interaction patterns, and then prune it to a suitable task-specific topology. Given the above limitations, a natural question arises: Going beyond template graph modification, can we design a more flexible and extensible paradigm for collaboration topology construction?
@@
MAS Collaboration Protocol. Given a collaboration graph $\mathcal{G}$, the MAS addresses a user query $\mathcal{Q}$ by executing a multi-step collaboration protocol. This protocol governs how information is processed and passed between agents, unfolding over a series of communication rounds. The operational sequence for agent activation within each round is determined by a topological sort of the nodes, ensuring that an agent is activated only after its prerequisite inputs are available. This process can be executed for $K$ rounds to allow for iterative refinement. In each round $k\in\{1,\dots,K\}$, an agent $v_{i}$ generates its response $m_{i}^{(k)}$ by invoking its language model with a dynamically constructed prompt $\mathcal{P}_{i}^{(k)}$:
@@
where the prompt integrates the intrinsic properties of the agent with the outputs of its predecessors from the previous round:
@@
After $K$ rounds, the final output $\mathcal{O}$ is obtained by aggregating the final-round responses from a subset of or all agents:
@@
[…]
@@
$P(\mathcal{G}|\mathcal{Q},\mathcal{R})=\prod_{i=1}^{|\mathcal{V}|}\underbrace{P(v_{i}|\mathcal{G}_{<i},\mathcal{Q},\mathcal{R})}_{\text{Node Generation}}\cdot\underbrace{\prod_{j=1}^{i-1}P(e_{ji}|v_{i},\mathcal{G}_{<i},\mathcal{Q})}_{\text{Edge Generation}},$ (5)
where $\mathcal{G}_{<i}$ represents the subgraph of the first $i-1$ nodes. The generation process at each step $i$ thus involves two key actions: node generation, predicting the role of the next agent to add, and edge generation, establishing its connections from existing agents. This formulation provides significant flexibility, enabling the model to dynamically determine the total number of agents by learning to sample a special END token, and to model complex structural dependencies by conditioning on the generation history.
@@
[…]
Method Task-Adaptive Variable Size Flexible Roles
[…]
@@
where $\sigma$ denotes the sigmoid function. This context, along with an edge feature vector $\mathbf{f}^{(i)}_{\text{edge}}$ (i.e., a vector encoding the connectivity pattern of the previously added node $v_{i-1}$), are concatenated into an input vector $[\mathbf{f}^{(i)}_{\text{cont}},\mathbf{f}^{(i)}_{\text{edge}}]$. Then, another GRU module, $\text{GRU}_{\text{node}}$, updates its hidden state $\mathbf{h}^{(i)}_{\text{node}}$, which captures the full generation condition:
@@
[…]
@@
Data Construction of Curriculum Learning.
$\mathcal{D}_{\text{eff}}=\mathcal{D}_{\text{simple}}\cup\mathcal{D}_{\text{pruned}}\cup\mathcal{D}_{\text{replay}}.$ (17)
Model Training.
@@
$\mathcal{L}_{\text{node}}$ $=-\sum_{(\mathcal{G},\mathcal{Q})\in\mathcal{D}}\sum_{i=1}^{|\mathcal{V}|}\log P_{\theta}(v_{i}|\mathcal{G}_{<i},\mathcal{Q}),$ (20)
@@
Table 2: Dataset descriptions and statistics.
@@
ARG-Designer 89.54 $\uparrow$9.15 94.37 $\uparrow$12.07 86.45 $\uparrow$15.39 98.93 $\uparrow$5.84 95.63 $\uparrow$9.08 91.74 $\uparrow$20.35 92.78 $\uparrow$11.98
Inference.
During inference, given a new task query $\mathcal{Q}$, the trained ARG-Designer model generates a graph autoregressively without ground-truth guidance. The process, detailed in Algo. 1, starts with the encoded task embedding and enters a generation loop. It samples a new agent role at each step, and if it is not the special END token, it then sequentially samples the existence of incoming edges from all previously generated nodes. This loop continues until an END token is sampled or a maximum node count is reached.
@@
Datasets and Metrics. Following (Zhang et al. 2025b), we evaluated ARG-Designer on three categories of datasets: ❶ General Reasoning: MMLU (Hendrycks et al. 2020); ❷ Mathematical Reasoning: GSM8K (Cobbe et al. 2021), MultiArith (Roy and Roth 2016), SVAMP (Patel, Bhattamishra, and Goyal 2021), and AQuA (Ling et al. 2017); ❸ Code Generation: HumanEval (Chen et al. 2021). The statistics of the datasets are shown in Table 2.
Baselines. We compare ARG-Designer against various baselines, which can be grouped into four main categories: ❶ Single-agent methods, including CoT (Wei et al. 2022) and Self-Consistency (Wang et al. 2022); ❷ MAS with fixed topologies, such as Chain, Tree, Complete Graph, and Random Graph (Qian et al. 2024); ❸ MAS with Debate like LLM-Debate (Du et al. 2023), where multiple agents iteratively critique and refine responses in a structured process; ❹ MAS with Learnable topologies, which include AgentPrune (Zhang et al. 2025a), AgentDropout (Wang et al. 2025), and G-Designer (Zhang et al. 2025b).
Implementation Details. We access GPT models via the OpenAI API, primarily using gpt-4o-2024-08-06 (GPT-4o). We employ a summarizer agent to aggregate the history of dialogue and produce the final solution $a^{(K)}$, with $K=3$ for all baselines across all experiments. The node encoder is implemented using all-MiniLM-L6-v2 (Wang et al. 2020), with the embedding dimension set to $D=384$. The hyperparameter $\alpha$ is set to 0.20.20.2 for all experiments. Following classical configurations in LLM-MAS (Zhuge et al. 2024; Yin et al. 2023; Zhang et al. 2025b), we provide explicit agent profiles for multi-agent methods and use GPT-4 to generate these profile pools. For all datasets, we use $B\in\{40,60\}$ queries for model training.
@@
Token Efficiency. A key benefit of ARG-Designer is its ability to generate tailored topologies for different tasks, which prevents unnecessary complexity and thus minimizes token consumption. Fig. 3a and 3b illustrate the trade-off between performance and token cost. We can observe that: ❶ ARG-Designer elegantly balances efficiency and performance. On the GSM8K dataset, ARG-Designer is the most token-efficient method, using only 4.1e6 tokens, while achieving a top-tier accuracy of 94.37%. It surpasses the strong G-Designer baseline in accuracy while using approximately 50% fewer tokens. ❷ A general trend where more complex communication structures, such as the dialogue-based LLM-Debate, achieve relatively good performance but at an extremely high token cost. ❸ A direct comparison between ARG-Designer and its w/o fine-tune variant further underscores the value of the efficiency fine-tuning phase. For instance, on MMLU, fine-tuning improves accuracy from 88.23% to 89.54% while simultaneously cutting token usage by nearly 30%. On GSM8K, it reduces token consumption by a massive 34% (from 6.25e6 to 4.1e6). This demonstrates that our two-phase training strategy is highly effective at optimizing for both performance and efficiency.
Ablation Study. To validate the effectiveness of key components in ARG-Designer, we investigate three major variants of ARG-Designer: ❶ w/o fine-tune, where efficiency fine-tuning phase is removed. ❷ w/o task embedding, where the influence of task embedding $\mathbf{f}_{\mathcal{Q}}$ is set to be 0. ❸ w/o hist. embedding, where the historical embeddings are removed from the generation model.
Table 4: Results of ablation study.
@@
ARG-Designer 89.54 94.37 91.74 91.88
w/o fine-tune 88.23 94.68 90.91 91.27
w/o task embedding 86.93 93.12 89.26 89.77
w/o hist. embedding 88.23 93.59 90.08 90.63
The results in Table 4 demonstrate the contribution of each component: ❶ ARG-Designer model achieves the highest average score of 91.88, significantly outperforming a vanilla baseline (78.02) that lacks these sophisticated mechanisms. ❷ The absence of task-specific guidance (w/o task embedding) causes the most significant performance degradation, which highlights that conditioning on the task is crucial for generating a bespoke and effective collaboration topology. ❸ The removal of historical embeddings (w/o hist. embedding) also leads to a noticeable decline to 90.63, confirming the value of modeling dependencies between agents. ❹ Interestingly, the performance of the w/o fine-tune variant is highly competitive. This indicates that the model can learn fundamental collaboration patterns from the first stage. Nevertheless, the fine-tuning can bring significant benefits in terms of efficiency (see Fig. 3), while also slightly improving the performance.
Robustness Analysis. Following Zhuge et al. (2024), we evaluate the robustness of ARG-Designer by simulating a system prompt attack, where an adversarial prompt is injected into a single agent to disrupt its function. As illustrated in Fig. 4a, this attack causes significant performance degradation in MAS with fixed and naive topologies. In contrast, ARG-Designer shows excellent robustness against attacks with the least performance degradation (2.15%). This resilience emerges from our training objective, which discourages brittle structures and guides the model to construct fault-tolerant topologies with distributed risk and redundant communication paths.
@@
Case Study. To further illustrate the advantages of ARG-Designer over learning-based baselines (e.g., G-Designer (Zhang et al. 2024)), we conduct a comparative case study on representative cases in HumanEval and GSM8K. As shown in Fig. 4c. The key difference lies in the flexibility of composition versus the static design. G‑Designer requires a predefined and fixed set of agents and a fixed agent count. Its solution graphs remain within this rigid template, regardless of task complexity. ARG-Designer, in contrast, dynamically generates both the roles and their communication links from an extensible role pool. It adapts the number of agents and connections based on task needs. Therefore, ARG-Designer constructs more efficient collaboration graphs with fewer agents and messages, cutting token usage without sacrificing accuracy.
@@
Autoregressive models are a cornerstone of graph generation, tackling the intractable joint probability of a graph by factorizing it into a tractable sequence of conditional probabilities for nodes and edges (You et al. 2018; Liao et al. 2019). The design of these models, however, is critically dependent on the choice of node ordering. Initial models like GraphRNN (You et al. 2018) pioneered this sequential approach but relied on simple, fixed node orderings like Breadth-First Search (BFS). Subsequent work like GRAN (Liao et al. 2019) showed that exploring various ad-hoc ordering schemes could improve performance, but highlighted the theoretical challenge that these orderings can lead to a loose variational bound. To resolve this, GraphGEN (Goyal, Jain, and Ranu 2020) proposed using a single, canonical order for a graph, though this creates a new mismatch problem as the generation process itself may not follow this canonical path. Beyond the pivotal node-ordering problem, other research has focused on improving scalability, with methods like BiGG (Dai et al. 2020) reducing generation time complexity and others proposing more scalable graph representations (Jang, Lee, and Ahn 2023). Another key direction is conditional generation, where models like CCGG (Ommi et al. 2022) learn to produce graphs based on a given class label, and hybrid architectures have also been explored, for instance by combining autoregressive models with diffusion models (Kong et al. 2023).
@@
The advent of powerful Large Language Models (LLMs) has catalyzed a shift from single-agent systems to Multi-Agent Systems (MAS) for tackling complex problems (Qian et al. 2024; Zhu et al. 2024). The success of these systems hinges on their collaboration topology, the design of which has thus become a central research problem (Zhuge et al. 2024; Zhang et al. 2024).
Research into topology design has evolved from static to adaptive structures. Initial approaches adopted fixed, manually-designed topologies, such as chains that enforce a sequential workflow (Wei et al. 2022; Hong et al. 2023) or trees that facilitate structured exploration and deliberation (Yao et al. 2023). While foundational, the inherent rigidity of these static structures limits their adaptability across diverse tasks, often leading to sub-optimal performance. To address this, a recent line of work has focused on learning adaptive communication graphs. For instance, some approaches start with a dense, fully-connected graph and learn to prune it in a task-aware manner. AgentPrune (Zhang et al. 2025a) learns to remove redundant communication links, while AgentDropout (Wang et al. 2025) applies a dynamic dropout technique to both agents and edges. To move beyond simple pruning, more advanced methods leverage the expressive power of graph neural networks (GNNs), which have become a standard for learning on graph-structured data (Li et al. 2024b; Chen et al. 2025). G-Designer (Zhang et al. 2025b), for instance, employs a GNN-based autoencoder to directly generate a query-dependent communication structure. While these methods achieve task-adaptivity in structuring communication, they are still fundamentally constrained by the initial template. This rigidity prevents the MAS from being truly bespoke, leading to the critical issues of redundant composition and limited extensibility.
Drawing inspiration from real-world practices of recruiting teams incrementally, our work departs fundamentally from modifying a predefined template. We instead propose autoregressive graph generation, a paradigm that constructs the collaboration graph from scratch. This allows our method, ARG-Designer, to jointly determine both the system’s composition and structure, dynamically selecting agents from an extensible pool and establishing optimal communication links in a truly task-adaptive manner.

### v4

@@ Abstract
Multi-agent systems (MAS) based on large language models (LLMs) have emerged as a powerful solution for dealing with complex problems across diverse domains. The effectiveness of MAS is critically dependent on its collaboration topology, which has become a focal point for automated design research. However, existing approaches are fundamentally constrained by their reliance on a template graph modification paradigm with a predefined set of agents and hard-coded interaction structures, significantly limiting their adaptability to task-specific requirements. To address these limitations, we reframe MAS design as a conditional autoregressive graph generation task, where both the system composition and structure are designed jointly. We propose ARG-Designer, a novel autoregressive model that operationalizes this paradigm by constructing the collaboration graph from scratch. Conditioned on a natural language task query, ARG-Designer sequentially and dynamically determines the required number of agents, selects their appropriate roles from an extensible pool, and establishes the optimal communication links between them. This generative approach creates a customized topology in a flexible and extensible manner, precisely tailored to the unique demands of different tasks. Extensive experiments across six diverse benchmarks demonstrate that ARG-Designer not only achieves state-of-the-art performance but also enjoys significantly greater token efficiency and enhanced extensibility.
Codehttps://github.com/Shiy-Li/ARG-Designer
@@ Introduction
Agents built on large language models (LLMs) have demonstrated impressive capabilities in tackling complex tasks across domains, including code generation, data analysis, decision-making, and question answering (zhu2024large; li2024dawn; song2023llm; wang2024chain; zhong2024debug; tan2025bisecle; miao2025blindguard). To overcome the limitations of a single agent in tackling more complex tasks, the research interests have increasingly shifted towards multi-agent systems (MAS), which unlock new potential through collaborative interactions among agents with diverse capabilities and roles. Central to MAS is its collaboration topology, a graph that defines how agents with roles are structured and how they exchange information (liu2025graph). A growing body of evidence, spanning from sequential reasoning pipelines to debate-based approaches, demonstrates that MAS performance varies dramatically depending on how inter-agent communication is architected (zhang2024aflow; zhou2025multi). Therefore, designing an effective collaboration graph tailored to specific tasks becomes a critical research challenge.
@@
Early research on MAS topology design focused on static and manually designed graphs, such as chains that enforce sequential workflow (wei2022chain; hong2023metagpt), trees that enable structured deliberation (yao2023tree), and fully connected graphs that ensure sufficient communication. Although these canonical collaboration topologies can facilitate effective coordination in specific scenarios, the inherent rigidity of these fixed topologies limits their adaptability across diverse tasks, resulting in sub-optimal performance. To enhance flexibility and efficiency, a more recent line of work (zhuge2024gptswarm; shen2025understanding) focuses on adaptively constructing task-specific communication structures using graph learning models (pan2025survey; pan2025label; fu2025less; fu2025learn). For example, AgentPrune (zhang2024cut) and AgentDrop (wang2025agentdropout) learn to create sparse and task-specific graphs by pruning connections or agents from a predefined template topology. More advanced approaches like G-Designer (zhang2024g) follow the paradigm of graph structure learning, which uses a graph autoencoder to learn efficient collaboration structures in a task-adaptive manner.
Despite their varied designs, existing graph learning-based methods often follow a shared paradigm: template graph modification (Fig. 1a). That is, they typically start from a fixed communication template based on a predefined set of agents and hard-coded interaction structures, and apply learnable adjustments, such as edge reweighting or pruning, to adapt the topology to specific tasks (zhang2024cut; zhang2024g). Despite offering reasonable adaptability in constrained settings, this paradigm exhibits two inherent limitations. Limitation 1: Redundant Composition. To ensure structural flexibility, template graphs are often initialized with numerous agent roles and densely connected edges, many of which are unnecessary for a specific task. Even with pruning mechanisms, irrelevant agents or connections may be retained in the learned task-specific topology, leading not only to reduced efficiency but also to potential sub-optimal decision-making during execution. Limitation 2: Limited Extensibility. In the fast-evolving field of LLM-based agents, a massive number of new agent functionalities are emerging with increasing frequency. However, trained on a fixed template graph, the existing methods struggle to generalize to scenarios with dynamic agent sets or evolving collaboration needs. Meanwhile, it would be prohibitively expensive to build a large-scale template graph that covers all possible agent roles and interaction patterns, and then prune it to a suitable task-specific topology. Given the above limitations, a natural question arises: Going beyond template graph modification, can we design a more flexible and extensible paradigm for collaboration topology construction?
@@
In this section, we introduce the graph-based modeling of MAS, and then formulate MAS topology design as an autoregressive graph generation problem.
@@
MAS Collaboration Protocol. Given a collaboration graph $\mathcal{G}$, the MAS addresses a user query $\mathcal{Q}$ by executing a multi-step collaboration protocol. This protocol governs how information is processed and passed between agents, unfolding over a series of communication rounds (pan2026correcting; zhang2024trustworthy; cai2024fgad). Unlike traditional GNNs message passing, the operational sequence for agent activation within each round is determined by a topological sort of the nodes, ensuring that an agent is activated only after its prerequisite inputs are available (zhao2025freegad; chen2025uncertainty; zhuang2025refine; li2024noise; cai2025agdiff). This process can be executed for $K$ rounds to allow for iterative refinement. In each round $k\in\{1,\dots,K\}$, an agent $v_{i}$ generates its response $m_{i}^{(k)}$ by invoking its language model with a dynamically constructed prompt $\mathcal{P}_{i}^{(k)}$:
@@
where the prompt integrates the properties of the agent with outputs of its predecessors from the previous round:
@@
After $K$ rounds, the final output $\mathcal{O}$ is obtained by aggregating the final-round responses from some or all agents:
@@
where the aggregation strategy $\text{Aggregate}(\cdot)$ varies across implementations. Common strategies include majority voting, delegating the final decision to a specific terminal agent, or selecting the output from the last agent in the execution order. The number of communication rounds $K$ can be either predefined or adaptively determined via early-stopping.
@@
[…]
where $\mathcal{G}_{<i}$ represents the subgraph of the first i1i-1 nodes. The generation process at each step $i$ thus involves two key actions: node generation, predicting the role of the next agent to add, and edge generation, establishing its connections from existing agents. This formulation provides significant flexibility, enabling the model to dynamically determine the total number of agents by learning to sample a special END token, and to model complex structural dependencies by conditioning on the generation history.
@@
Method Task- Variable Flexible
Adaptive Size Roles
Manual Design
AgentDropout
AgentPrune
G-Designer
ARG-Designer (ours)
Table 1: Comparison of degrees of freedom in MAS design paradigms. The icons ✓, –, and ✗ represent full, partial, and no support for each capability, respectively.
@@
where $\sigma$ denotes the sigmoid function. This context, along with an edge feature vector $\mathbf{f}^{(i)}_{\text{edge}}$ (i.e., a vector encoding the connectivity pattern of the previously added node $v_{i-1}$), are concatenated into an input vector $\mathbf{f}^{(i)}_{\text{cont}},\mathbf{f}^{(i)}_{\text{edge}}$]. Then, another GRU module, $\text{GRU}_{\text{node}}$, updates its hidden state $\mathbf{h}^{(i)}_{\text{node}}$, which captures the full generation condition:
@@
where the $\text{Softmax}(\cdot)$ function converts the scores into a probability distribution.
@@
Data Construction of Curriculum Learning.
$\mathcal{D}_{\text{eff}}=\mathcal{D}_{\text{simple}}\cup\mathcal{D}_{\text{pruned}}\cup\mathcal{D}_{\text{replay}},$ (17)
Model Training.
@@
$\mathcal{L}_{\text{node}}$ $=-\sum_{(\mathcal{G},\mathcal{Q})\in\mathcal{D}}\sum_{i=1}^{|\mathcal{V}|}\log P_{\theta}(v_{i}|\mathcal{G}_{<i},\mathcal{Q},\mathcal{R}),$ (20)
@@
Inference.
During inference, given a new task query $\mathcal{Q}$, the trained ARG-Designer model generates a collaboration graph autoregressively without any ground-truth guidance. The process begins by initializing an empty graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. It then enters a generation loop that iteratively builds the graph node by node. At each step $i$, the model’s node generator first samples a role $r_{i}$ for the new agent $v_{i}$ from the probability distribution $P_{\theta}(v_{i}|\mathcal{G}_{<i},\mathcal{Q})$. If the sampled role is the special END token, or if a predefined maximum number of agents $N_{\max}$ is reached, the generation process terminates. Otherwise, the new node $v_{i}$ is added to the vertex set $\mathcal{V}$. Subsequently, the model’s edge generator is invoked. It sequentially considers each existing node $v_{j}\in\mathcal{G}_{<i}$ and samples the existence of an incoming edge $e_{ji}$ from $v_{j}$ to $v_{i}$. This entire process of node and edge generation is repeated until a termination condition is met, at which point the final graph $\mathcal{G}$ is returned. A detailed algorithmic description see Algo. 1.
@@
ARG-Designer 89.54 $\uparrow$9.15 94.40 $\uparrow$12.10 86.45 $\uparrow$15.39 98.93 $\uparrow$5.84 95.63 $\uparrow$9.08 91.74 $\uparrow$20.35 92.78 $\uparrow$11.98
@@
Datasets and Metrics. Following (zhang2024g), we evaluated ARG-Designer on three categories of datasets: ❶ General Reasoning: MMLU (hendrycks2020measuring); ❷ Mathematical Reasoning: GSM8K (cobbe2021training), MultiArith (roy2016solving), SVAMP (patel2021nlp), and AQuA (ling2017program); ❸ Code Generation: HumanEval (chen2021evaluating). The statistics of the datasets are shown in Table 2.
Baselines. We compare ARG-Designer against various baselines, which can be grouped into four main categories: ❶ Single-agent methods, including CoT (wei2022chain) and Self-Consistency (wang2022self); ❷ MAS with fixed topologies, such as Chain, Tree, Complete Graph, and Random Graph (qian2024scaling); ❸ MAS with Debate like LLM-Debate (du2023improving), where multiple agents iteratively critique and refine responses in a structured process; ❹ MAS with Learnable topologies, which include AgentPrune (zhang2024cut), AgentDropout (wang2025agentdropout), and G-Designer (zhang2024g).
Implementation Details. We access GPT models via the OpenAI API, primarily using gpt-4o-2024-08-06 (GPT-4o). We employ a summarizer agent to aggregate the history of dialogue and produce the final solution $a^{(K)}$, with $K=3$ for all baselines across all experiments. The node encoder is implemented using all-MiniLM-L6-v2 (wang2020minilm), with the embedding dimension set to $D=384$. The hyperparameter $\alpha$ is set to $0.2$ for all experiments. Following classical configurations in LLM-MAS (zhuge2024gptswarm; yin2023exchange; zhang2024g), we provide explicit agent profiles for multi-agent methods and use GPT-4 to generate these profile pools. For all datasets, we use $B\in\{40,60\}$ queries for model training.
@@
Token Efficiency. A key benefit of ARG-Designer is its ability to generate tailored topologies for different tasks, which prevents unnecessary complexity and thus minimizes token consumption. Fig. 3a and 3b illustrate the trade-off between performance and token cost. We can observe that: ❶ ARG-Designer elegantly balances efficiency and performance. On the GSM8K dataset, ARG-Designer is the most token-efficient method, using only 4.1e6 tokens, while achieving a top-tier accuracy of 94.40%. It surpasses the strong G-Designer baseline in accuracy while using approximately 50% fewer tokens. ❷ A general trend where more complex communication structures, such as the dialogue-based LLM-Debate, achieve relatively good performance but at an extremely high token cost. ❸ A direct comparison between ARG-Designer and its w/o fine-tune variant further underscores the value of the efficiency fine-tuning phase. For instance, on MMLU, fine-tuning improves accuracy from 88.23% to 89.54% while simultaneously cutting token usage by nearly 30%. On GSM8K, it reduces token consumption by a massive 34% (from 6.25e6 to 4.1e6). This demonstrates that our two-phase training strategy is highly effective at optimizing for both performance and efficiency.
@@
Ablation Study. To validate the effectiveness of key components in ARG-Designer, we investigate three major variants of our model: ❶ w/o fine-tune, where efficiency fine-tuning phase is removed. ❷ w/o task emb., where the influence of task embedding $\mathbf{f}_{\mathcal{Q}}$ is set to be 0. ❸ w/o hist. emb., where the historical embeddings are removed from the generation model. The results in Table 4 demonstrate the contribution of each component: ❶ ARG-Designer achieves the highest average score of 91.89, significantly outperforming a vanilla baseline (78.02) that lacks these sophisticated mechanisms. ❷ The absence of task-specific guidance (w/o task emb.) causes the most significant performance degradation, which highlights that conditioning on the task is crucial for generating a bespoke and effective collaboration topology. ❸ The removal of historical embeddings (w/o hist. emb.) also leads to a noticeable decline to 90.64, confirming the value of modeling dependencies between agents. ❹ Interestingly, the performance of the w/o fine-tune variant is highly competitive. This indicates that the model can learn fundamental collaboration patterns from the first stage. Nevertheless, the fine-tuning bring significant efficiency benefits (see Fig. 3), while also slightly improving performance.
@@
ARG-Designer 89.54 94.40 91.74 91.89
w/o fine-tune 88.23 94.70 90.91 91.28
w/o task emb. 86.93 93.10 89.26 89.76
w/o hist. emb. 88.23 93.60 90.08 90.64
Table 4: Results of ablation study.
Robustness Analysis. Following zhuge2024gptswarm, we evaluate the robustness of ARG-Designer by simulating a system prompt attack, where an adversarial prompt is injected into a single agent to disrupt its function. As illustrated in Fig. 4a, this attack causes significant performance degradation in MAS with fixed and naive topologies. In contrast, ARG-Designer shows excellent robustness against attacks with the least performance degradation (2.15%). This resilience emerges from our training objective, which discourages brittle structures and guides the model to construct fault-tolerant topologies with distributed risk and redundant communication paths.
@@
Case Study. To further illustrate the advantages of ARG-Designer over learning-based baselines (e.g., G-Designer (zhang2024aflow)), we conduct a comparative case study on representative cases in HumanEval and GSM8K. As shown in Fig. 4c. The key difference lies in the flexibility of composition versus the static design. G‑Designer requires a predefined and fixed set of agents and a fixed agent count. Its solution graphs remain within this rigid template, regardless of task complexity. ARG-Designer, in contrast, dynamically generates both the roles and their communication links from an extensible role pool. It adapts the number of agents and connections based on task needs. Therefore, ARG-Designer constructs more efficient collaboration graphs with fewer agents and messages, cutting token usage without sacrificing accuracy.
@@
Autoregressive models are the cornerstone of graph generation, tackling the intractable joint probability of a graph by factorizing it into a tractable sequence of conditional probabilities for nodes and edges (you2018graphrnn; liao2019efficient). However, the design of these models is critically dependent on the choice of node ordering. Initial models like GraphRNN (you2018graphrnn) pioneered this sequential approach but relied on simple, fixed node orderings like Breadth-First Search (BFS). Subsequent work such as GRAN (liao2019efficient) showed that exploring various ad-hoc ordering schemes could improve performance, but highlighted the theoretical challenge that these orderings can lead to a loose variational bound. To resolve this, GraphGEN (goyal2020graphgen) proposed using a single canonical order for a graph, although this creates a new mismatch problem, as the generation process itself may not follow this canonical path. Beyond the pivotal node-ordering problem, other research has focused on improving scalability, with methods like BiGG (dai2020scalable) reducing generation time complexity and others proposing more scalable graph representations (jang2023simple). Another key direction is conditional generation, where models like CCGG (ommi2022ccgg) learn to produce graphs based on a given class label, and hybrid architectures have also been explored, for instance by combining autoregressive models with diffusion models (kong2023autoregressive).
@@
The advent of powerful Large Language Models (LLMs) has catalyzed a shift from single-agent systems to Multi-Agent Systems (MAS) for tackling complex problems (qian2024scaling; zhu2024large). The success of these systems hinges on their collaboration topology, the design of which has thus become a central research problem (zhuge2024gptswarm; zhang2024aflow).
Research into topology design has evolved from static to adaptive structures. The initial approaches adopted fixed and manually-designed topologies, such as chains that enforce a sequential workflow (wei2022chain; hong2023metagpt) or trees that facilitate structured exploration and deliberation (yao2023tree). While foundational, the inherent rigidity of these static structures limits their adaptability across diverse tasks, often leading to sub-optimal performance. To address this, a recent line of work has focused on learning adaptive communication graphs. For instance, some approaches start with a dense, fully-connected graph and learn to prune it in a task-aware manner. AgentPrune (zhang2024cut) learns to remove redundant communication links, while AgentDropout (wang2025agentdropout) applies a dynamic dropout technique to both agents and edges. To move beyond simple pruning, more advanced methods leverage the expressive power of graph neural networks (GNNs), which have become a standard for learning on graph-structured data (li2024noise; chen2025uncertainty). G-Designer (zhang2024g), for instance, employs a GNN-based autoencoder to directly generate a query-dependent communication structure. While these methods achieve task-adaptivity in structuring communication, they are still fundamentally constrained by the initial template. This rigidity prevents the MAS from being truly bespoke, leading to the critical issues of redundant composition and limited extensibility.
Drawing inspiration from the real-world practices of recruiting teams incrementally, our work departs fundamentally from modifying a predefined template. We instead propose autoregressive graph generation, a paradigm that constructs the collaboration graph from scratch. This allows our method ARG-Designer, to jointly determine both the system’s composition and structure, dynamically selecting agents from an extensible pool and establishing optimal communication links in a truly task-adaptive manner.